OpenStudy (tiffany_rhodes):

Find the directional derivative of the function f at the point p in the direction of the vector v, where f(x,y)= e^(xy) * (cosx+siny); p=(pi/2,0); v=5i-j.

OpenStudy (tiffany_rhodes):

I know that to find the directional derivative, you find the gradient of f. So you compute the partial derivatives of f(x,y) wrt x and y. Then you take the dot product of the gradient of f with the vector v.

OpenStudy (tiffany_rhodes):

I'm confused about the vector v. My homework won't accept the result of the dot product in terms of i,j, or k.

OpenStudy (anonymous):

they probably just mean \[\nabla f(x,y) \cdot \vec{v} =\frac{\partial f}{\partial x}\hat{i} \cdot 5\hat{i}-\frac{\partial f}{\partial y}\hat{j} \cdot \hat{j} + 0\hat{k} \cdot \hat{k} = etc\] and then just jour answer in the end

OpenStudy (dan815):

not wth v, unitvector v

OpenStudy (dan815):

or else u are getting a project of your gradient on that vector * the magnitude of that vector

OpenStudy (dan815):

projection*

OpenStudy (anonymous):

that depends, have a look at http://en.wikipedia.org/wiki/Directional_derivative#Generally_applicable_definition I don't know the context of that homework, so I went for the most general definition, that tiffany_rhodes also mentioned in the second part here Anyway, that was not the point, the point was, where do the terms i, j and k come in. You only see them if you write it out like in components It can also be done in a way that you do not write i j and ka anywhaere:

OpenStudy (anonymous):

like this \[\nabla f(x,y) \cdot \vec{v} = \begin{pmatrix} \frac{\partial f}{\partial x}\\ \frac{\partial f}{\partial y}\\ 0 \end{pmatrix} \cdot \begin{pmatrix} 5\\ -1\\ 0 \end{pmatrix}\]

OpenStudy (anonymous):

Then you forget about the i j and k

OpenStudy (anonymous):

and it is also true that they drop out of the final answer